# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ 2.2 Incidental Parameter Problem
$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\big{(}\widehat{\psi}\left(\theta\right)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)},$
with $a=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1),$ and $\psi_{j}$ denotes the $j$th component of $\psi.$ Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j},$ evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right).$ For notational simplicity, we write
$s(\theta):=s\left(\theta,\psi\left(\theta\right)\right)\quad\text{and}\quad H\left(\theta\right):=H\left(\theta,\psi(\theta)\right).$ (2.4)
From (2.3), by Assumption 1, we obtain that
$\widehat{\psi}(\theta)-\psi(\theta)$ $=[-{H}(\theta)]^{-1}\big{(}s(\theta)+r^{s}\left(\theta,a\right)\big{)}.$ (2.5)
Next, define the profile likelihoods as $\widehat{l}(\theta):=l^{*}(\theta,\psi)\mid_{\psi=\widehat{\psi}(\theta)}$ and $l(\theta):=l^{*}(\theta,\psi)\mid_{\psi={\psi}(\theta)}$. A Taylor expansion of $\widehat{l}(\theta)$ around $\psi(\theta)$ gives
$\widehat{l}(\theta)=l(\theta)+s^{\prime}\left(\theta\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}{H}(\theta)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{l}\left(\theta,a_{0}\right)$ (2.6)
where $r^{l}\left(\theta,a_{0}\right):=1/3\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{(N+T-2)K})$ for some $a_{0}\in(0,1).$ Combining (2.5) and (2.6), we have
$\widehat{l}(\theta)=l(\theta)-\frac{1}{2}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)+r(\theta)$ (2.7)
where $r(\theta):=1/2(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)+r^{l}\left(\theta,a_{0}\right).$ Here we suppress $a$ and $a_{0}$ in $r(\theta)$ as they are constants. Under mild regularity conditions, we show that
$\sqrt{NT}\mathbb{E}\big{\{}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\big{\}}\nrightarrow 0$
and that $r(\theta)$ is a remainder term that is negligible for inference (see Section 3). This implies that the presence of $s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)$ in (2.7), due to the estimation of $\psi(\theta),$ contributes the MLE of $\theta_{0}$ based on $\widehat{l}(\theta).$ Consequently, the structural estimators and $\widehat{l}(\theta)$ based test statistics no longer follow the standard asymptotic distributions as in classical maximum likelihood theory. This is referred to as the IPP in this context. We use the following example to illustrate IPP.
Example 2 (heterogeneous means)
Assume that $Y_{it}$ follows a normal distribution with mean $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and unit variance, independent across both $i$ and $t.$ $\theta_{0}$ is the structural parameter of interest, while $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ are fixed effects. The likelihood function is
$l\left(\theta,\phi\right)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}-\frac{1}{2}\left(Y_{it}-\theta-\alpha_{i}-\gamma_{t}\right)^{2}-\frac{1}{2}\log(2\pi).$
Define $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}.$ Under the identification condition, the reparameterized likelihood $l^{\ast}(\theta,\psi)=l(\theta,D^{\prime}\psi).$ Given $\theta,$ maximizing $l^{\ast}(\theta,\psi)$ with respect to $\psi$ yields
$\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad i=1,\ldots,N-1,$ (2.8)
$\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad t=1,\ldots,T-1.$ (2.9)
Let $\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta)\big{)}^{\prime}=:\widehat{\psi}(\theta)$ and $\big{(}{\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0}\big{)}^{\prime}=:\psi_{0}.$ Define $\widehat{\alpha}_{N}(\theta):=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}(\theta)$ and $\widehat{\gamma}_{T}(\theta):=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}(\theta).$ Note that (2.8) and (2.9) also hold for $i=N$ and $t=T$ respectively. Denote $\phi_{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\alpha}_{N}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0},{\gamma}_{T}^{0})^{\prime}$ and $\widehat{\phi}(\theta):=\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\alpha}_{N}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta),\widehat{\gamma}_{T}(\theta)\big{)}^{\prime}.$ Note that
$\widehat{l}(\theta)-l(\theta)$ $=$ $l(\theta,\widehat{\phi}(\theta))-l(\theta,\phi_{0})$
$=$ $\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})\big{]}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})$
$+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})\big{]}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})$
where $Y_{it}(\theta):=Y_{it}-\theta-\alpha_{i}^{0}-\gamma_{t}^{0}.$ Since $Y_{it}$ is independent across $i$ and $t,$ by substituting (2.8) and (2.9), we obtain
$\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}$ $=\frac{1}{T}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+\frac{1}{N}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+o\big{(}1/\sqrt{NT}\big{)}$ (2.10)
$=\frac{1}{2}\frac{1}{T}+\frac{1}{2}\frac{1}{N}+o\big{(}1/\sqrt{NT}\big{)}.$
This leads to the conclusion that $\sqrt{NT}\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}\nrightarrow 0,$ revealing the IPP.
2.3 Bias Correction
The expansion in (2.10) from Example 2 motivates the correction of the likelihood $\widehat{l}(\theta)$ by subtracting the first two terms on the right-hand side of (2.10), ensuring that the corrected likelihood converges to $l(\theta)$ at a rate faster than $1/\sqrt{NT}.$ This paper proposes an analytical bias correction procedure by estimating the bias term $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ from (2.7) under $N/T\rightarrow\kappa\in(0,\infty)$. Specifically, let $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta),$ where $\breve{\alpha}:=(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime})^{\prime}$ and $\breve{\gamma}:=(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime})^{\prime}.$ Accordingly, we partition $[\mathbb{E}H(\theta)]^{-1}$ as
$[\mathbb{E}H(\theta)]^{-1}=:\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
where $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)$ are $(N-1)K\times(N-1)K$ and $(T-1)K\times(T-1)K$ diagonal blocks, respectively, and $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)=[H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)]^{\prime}.$ Lemma 1 (see Section 3) shows that $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ can be asymptotically decomposed into two bias components: $1/2\mathbb{E}\{s_{\breve{\alpha}}^{\prime}(\theta)[H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]s_{\breve{\alpha}}(\theta)\}=:{B}_{\breve{\alpha}}(\theta)$ and $1/2\mathbb{E}\{s_{\breve{\gamma}}^{\prime}(\theta)[H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)]s_{\breve{\gamma}}(\theta)\}=:{B}_{\breve{\gamma}}(\theta),$ resulting from the estimation of the individual and time effects respectively. By estimating the two bias components, we construct the corrected likelihood as follows:
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta).$ (2.11)
The explicit forms of $\widehat{B}_{\breve{\alpha}}(\theta)$ and $\widehat{B}_{\breve{\gamma}}(\theta)$ are defined in Section 3. Subsequently, Theorem 2 demonstrates that
$L(\theta)-l(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
implying that the corrected likelihood no longer suffers from the IPP. Define the MLE of $\theta_{0}$ as the maximizer of $L(\theta):$
$\widehat{\theta}_{L}=\arg\max_{\theta\in\Theta}L(\theta),$ (2.12)
@@ 2.3 Bias Correction
where $\Theta$ is a compact subset of $\mathbb{R}^{K}.$ Under mild regularity conditions, this MLE is asymptotically unbiased, and its asymptotic properties align with the standard results for classical MLE in the independent and identically distributed (i.i.d.) case. Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis.
Since the corrected likelihood depends implicitly on $\widehat{\psi}(\theta),$ which maximizes the reparameterized likelihood function $l^{\ast}(\theta,\psi),$ obtaining $\widehat{\theta}_{L}$ from $L(\theta)$ involves a two-layer optimization. In the inner layer, $\widehat{\psi}(\theta)$ is found by maximizing $l^{\ast}(\theta,\psi)$ using the Newton-Raphson algorithm with analytical derivatives. In the outer layer, we maximize $L(\theta)$ with respect to $\theta,$ substituting $\widehat{\psi}(\theta)$ and using numerical derivatives.
3 Asymptotic Theory
In this section, we develop the asymptotic theory for our bias correction procedure, as well as for the estimation and testing based on the corrected likelihood function. Let $\phi(\theta):=D^{\prime}\psi(\theta)$ and $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta).$ The following assumptions are introduced.
@@ 3 Asymptotic Theory
  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N$ and $T\to\infty.$
  • (ii) Let $\Theta$ be a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior. For each $\theta\in\Theta,$ there exists a compact subset $\Lambda_{\theta}$ that contains both $\widehat{\phi}(\theta)$ and $\phi(\theta)$ in its interior, such that $l(\theta,\phi)$ is three-times continuously differentiable with respect to $\phi\in\Lambda_{\theta}$ almost surely. Let $w_{it}=(Y_{it},X_{it}^{\prime})^{\prime},$ and assume there exists a function $g(w_{it})>0$ independent of $\theta$ and $\phi,$ such that
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Lambda_{\theta}}\left|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ understood as no differentiation), and $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\left[g^{\eta}(w_{it})\right]<\infty,$ for some $\eta>2.$
  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{ik}=\sum_{t=1}^{T}\gamma_{tk}=0$.
  • (iv) Given any fixed $\theta\in\Theta,$ for all $\phi\in\Lambda_{\theta}$ if $\phi\neq\phi(\theta),$ $\mathbb{P}\big{(}l(\theta,\phi)\neq l(\theta,\phi(\theta))\big{)}>0.$
@@
  • (v) Conditional on $\phi(\theta),$$\left\{w_{it}\right\}$ are independent across $i$. Denote $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ to be the $\sigma$-algebras generated, respectively, by $\left\{w_{is}:1\leq s\leq t\right\}$ and $\left\{w_{is}:t\leq s\leq T\right\}$. $\left\{w_{it}\right\}$ is conditional strong mixing with mixing coefficient

### v4

@@ 2.2 Incidental Parameter Problem
$\mathbb{E[}l(\theta,\widehat{\psi})-l(\theta,\psi^{0})]=-B(\theta)+O(1/(NT)),\qquad B(\theta)=-\left(\frac{1}{T}+\frac{1}{N}\right)\frac{\sigma^{0}}{2\sigma}.$
The term $B(\theta)$ is the IPP bias. Apparently,
$\sqrt{NT}B(\theta)=-\left(\sqrt{\frac{N}{T}}+\sqrt{\frac{T}{N}}\right)\frac{\sigma^{0}}{2\sigma}\not\rightarrow 0$
as $N,T\rightarrow\infty$ and $N/T\rightarrow\kappa$. When $N/T\rightarrow 0$ (or $N/T\rightarrow\infty$), $\sqrt{T/N}\rightarrow\infty$ (or $\sqrt{N/T}\rightarrow\infty$) and $\sqrt{NT}B(\theta)$ explodes.
The bias term $B(\theta)$ causes the variance estimator to be biased. We illustrate this consequence here. Observe that
$l(\theta,\widehat{\psi})$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log\sigma-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2},$
$\partial_{\sigma}l(\theta,\widehat{\psi})$ $=-\frac{1}{2\sigma}+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma^{2}}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}.$
Following the same steps in Example 2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (2.10); and it is independent of $\sigma$. Therefore, equating $\partial_{\sigma}l(\theta,\widehat{\psi})=0$, substituting $\mu$ with $\widehat{\mu}$, and solving for $\sigma$, the estimator of $\sigma^{0}$ is
$\widehat{\sigma}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}(Y_{it}-\widehat{\mu}-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}$
$=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}[Y_{it}-\mu_{0}-\alpha_{i}^{0}-\gamma_{t}^{0}-(\widehat{\mu}-\mu_{0})-(\widehat{\alpha}_{i}-\alpha_{i}^{0})-(\widehat{\gamma}_{t}-\gamma_{t}^{0})]^{2},$
where it is not difficult to see
$\widehat{\mu}-\mu_{0}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$
Therefore, seeing $\widetilde{Y}_{it}=Y_{it}-\mu_{0}-\alpha_{i}^{0}-\gamma_{t}^{0}$ by definition, we finally obtain
$\widehat{\sigma}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}^{2}-\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}-\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
$\quad\quad+\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T},\widetilde{Y}_{it}\widetilde{Y}_{js}$
from where we conclude that
$\mathbb{E}\widehat{\sigma}=\sigma_{0}-\frac{\sigma_{0}}{T}-\frac{\sigma_{0}}{N}+\frac{\sigma_{0}}{NT},$
leading to
$\sqrt{NT}(\mathbb{E}\widehat{\sigma}-\sigma_{0})\rightarrow-\kappa\sigma_{0}-\frac{\sigma_{0}}{\kappa}$
as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$. While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $-\sigma_{0}/T$ and $-\sigma_{0}/N$. These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $-\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
3 Bias Correction and Asymptotic Theory
The bias expansion of $l(\theta,\widehat{\psi}(\theta))$ in Equation (2.14) and Example 3 above motivate our approach of correcting the likelihood $l(\theta,\widehat{\psi}(\theta))$ by adding $B(\theta)$. The corrected likelihood can be viewed as an approximation to the IPP-free infeasible likelihood $l(\theta,\psi(\theta))$, with approximation error $o(1/\sqrt{NT})$. In Theorem 1 below we formalize the expansion in Equation (2.14). Specifically, we show that $B(\theta)$ can be further decomposed into $B(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT})$, in which $B_{\alpha}(\theta)$ appears because of the estimation errors of the individual-specific heterogeneities $\alpha_{i}^{0}$, while $B_{\gamma}(\theta)$ is present due to the estimation errors of the time-specific heterogeneities $\gamma_{t}^{0}$. The exact formulas for both terms are given below. Theorem 2 follows to establish the corrected likelihood as
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)$ (3.1)
where we denote $\widehat{l}(\theta):=l(\theta,\widehat{\psi}(\theta))$ for simplicity from now on. $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are estimates of $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, respectively. The analytical formulas for $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are provided below. We then show that
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
where we use $l(\theta):=l(\theta,\psi(\theta))$ for simplicity below. This implies that $L(\theta)$ does not suffer from the IPP. The corrected estimator of $\theta_{0}$, and the LR, LM, and Wald test statistics, may subsequently be derived from $L\left(\theta\right)$, instead of $\widehat{l}(\theta)$. Specifically, the corrected estimator of $\theta_{0}$ is
$\widehat{\theta}_{L}:=\arg\max_{\theta}L(\theta).$
@@ 3 Bias Correction and Asymptotic Theory
Under mild regularity conditions, $\widehat{\theta}_{L}$ is asymptotically unbiased, and its asymptotic properties align with the standard results of the classical MLE without the nuisance parameters. Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis. These results are formally stated in Corollaries 1 and 2, respectively.
Remark 2 (computation)
The computation of $\widehat{\theta}_{L}$ involves a two-layer maximization but can be done efficiently. In the inner layer, $\widehat{\psi}(\theta)$ is obtained by maximizing $l(\theta,\psi)$, keeping $\theta$ fixed. This is large-scale maximization problem and a few strategies can be exploited to expedite the procedure:
  1. 1.
    For many commonly used models, the first and second derivatives of $l(\theta,\psi)$ have clear analytical expressions, so that the standard Newton-Raphson algorithm can be employed to obtain $\widehat{\psi}(\theta)$.
  2. 2.
    Many models (specifically, generalized linear models) are equipped with linear indices, as in Example 1. In that case, the chain-rule-based technique in Sun and Dhaene (2019) can be used to achieve a fast calculation of the derivatives. The computational complexity of such a technique only grows with the number of linear indices, not the dimension of $\psi$, even without analytical expressions for the derivatives of $l(\theta,\psi)$.
  3. 3.
    The derivatives of the reparameterized $l(\theta,\psi)$ can be identically constructed from the derivatives of the original likelihood $\ell(\theta,\phi)$. Specifically, the reparameterization implies
    $\partial_{\psi}l(\theta,\psi)=D\partial_{\phi}\ell(\theta,\phi),\qquad\partial_{\psi\psi^{\prime}}l(\theta,\psi)=D\partial_{\phi\phi^{\prime}}\ell(\theta,\phi)D^{\prime},$
    holding true for every $\phi=D^{\prime}\psi$ and every $\theta$.
In the outer layer where $L(\theta)$ is maximized w.r.t. $\theta$, it can be tedious to obtain analytical formulas for the derivatives of the corrected likelihood $L(\theta)$. However, the dimension of $\theta$ is usually small, so that invoking the Newton-Raphson algorithm with numerical derivatives generally does not suffer from too much computational pressure, provided that the inner-layer maximization is done efficiently.
Now, we proceed to our bias correction procedure on the likelihood and present the related asymptotic theory. Denote $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta)$ and $\phi(\theta):=D^{\prime}\psi(\theta)$, which are essentially the “unparameterized” counterparts of $\widehat{\psi}(\theta)$ and $\psi(\theta)$, respectively. We impose the following assumption.
@@
  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N,T\rightarrow\infty$.
[…]
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Phi}\left|\partial_{\phi_{1}\cdots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ is understood as not taking derivatives), and
    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\{[g(w_{it})]^{\eta}\}<\infty,$
    for some $\eta>2$.
  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{k,i}^{0}=\sum_{t=1}^{T}\gamma_{k,t}^{0}=0$, where $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ are the $k$-th component of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, respectively.
  • (iv)  For each $\theta\in\Theta$, $\mathbb{P}[l(\theta,\psi)\neq l(\theta,\psi(\theta))]>0$ for every $\psi$ such that $D^{\prime}\psi\in\Phi$ and $\psi\neq\psi(\theta)$.
@@
  • (v) Conditional on $\phi^{0}$, $\{w_{it}\}$ are independent across $i$ and conditionally strong mixing with mixing coefficient
    $a_{i}(m):=\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}(A\cap B|\phi^{0})-\mathbb{P}(A|\phi^{0})\mathbb{P}(B|\phi^{0})\right|$
    such that, for $\eta>2$ as above,
    $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}[a_{i}(m)]^{1-2/\eta}<\infty,$
    where $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ are the $\sigma$-algebras generated by $\{w_{is}:1\leq s\leq t\}$ and $\{w_{is}:t\leq s\leq T\}$, respectively.
  • (vi) The matrix has $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
    $\sup_{N>N_{0},T>T_{0}}\max_{1\leq p\leq(N+T-2)K}\sup_{\theta\in\Theta}\lambda_{p}(\theta)<0$
    for some integers $N_{0}$ and $T_{0}$.
Here Assumption 1 is the same to Fernández-Val and Weidner (2016). It is used to ensure that the bias $\sqrt{NT}B(\theta)$ does not explode as $N,T\rightarrow\infty$, as discussed above. Assumption 1 is standard in the literature. See, e.g., Hahn and Newey (2004), Hahn and Kuersteiner (2011), and Arellano and Hahn (2016). It simply requires that the likelihood function $l_{i,t}(\theta,\phi)$ is differentiable in $\phi$ at least three times. The term $B(\theta)$ in (2.14) depends on the first and second derivatives, and the remainder $r(\theta)$ depends on the third. We require $l_{i,t}(\theta,\phi)$ and its derivatives to be uniformly bounded by a function $g(w_{it})$ whose moment exists to an order greater than the second. This is used to ensure, amongst other things, i) the interchangeability of $\mathbb{E}$ and $\partial$; and ii) the existence of moments, such as $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ in (2.14), for every $\theta\in\Theta$. It is also used to establish the uniform rate of convergence, over $\theta\in\Theta$, of the corrected likelihood $L(\theta)$. The uniform convergence of $L(\theta)$ guarantees that the bias correction effect carries over to both the estimators and likelihood-based test statistics. Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 00 over $i=1,\ldots,N$ and $t=1,\ldots,T$. Given the additive structure of $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ in (2.1), this condition is without loss of generality because we merely let $\theta_{0}$ capture the average effects of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Assumption 1 is an identification condition for $\psi(\theta)$. It is standard for maximum likelihood models. See, e.g., Newey and McFadden (1994). Assumption 1 is common in the field of dynamic panels. See, for instance, Hahn and Kuersteiner (2007), Fernández-Val and Weidner (2016), and Dhaene and Jochmans (2015). Notice that, the argument is conditional on the true heterogeneous effects $\phi^{0}$. Without conditioning, $w_{it}$ may not be, for instance, independent over $i$, because $w_{it}$ and $w_{jt}$ may depend on the same $\gamma_{t}^{0}$. In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $[a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$. The mixing condition is used to guarantee that certain autocovariances which appear in our proof are summable by Davydov’s inequality. Notice that we are not requiring $w_{it}$ to be identically distributed over $i$ or stationary over $t$. These are generally not possible because $w_{it}$ depends on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 00 for all $N>N_{0}$ and $T>T_{0}$. This ensures that $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ is well defined for large $N$ and $T$. This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood. We state the condition with “less than 00” because the Hessian in a maximum likelihood model is usually negative definite.
Next, partition [^5] Note that $s(\theta)=s(\theta)-\mathbb{E}s\left(\theta\right)$ because $\mathbb{E}s\left(\theta\right)=0$ by definition. We keep $\mathbb{E}s\left(\theta\right)$ here for the convenience of constructing the corrected likelihood. $s(\theta)-\mathbb{E}s\left(\theta\right)$ and $[\mathbb{E}H(\theta)]^{-1}$ as